OpenStudy (anonymous):

Which of the following is the radical expression of 4d^3/8?

OpenStudy (lgbasallote):

hint \[\huge \sqrt[m]{x^n} \implies x^{n/m}\] does that help?

OpenStudy (anonymous):

no..im sorry i'm no good with this..

OpenStudy (anonymous):

waaaait!

OpenStudy (anonymous):

is it ^3/4d8

OpenStudy (lgbasallote):

okay...maybe a numerical example \[\huge a^{2/3} \implies \sqrt[3]{a^2}\] \[\huge 5a^{4/5} \implies 5\sqrt[5]{a^4}\] i think my second example cn tell you if you're right or wrong ;)

OpenStudy (lgbasallote):

\[\huge (5a)^{4/5} \implies \sqrt[5]{5a^4}\]

OpenStudy (lgbasallote):

so what do you think is the answer for \[\huge 4d^{3/8}\]

OpenStudy (anonymous):

damn I was waaay off, thnx tho

OpenStudy (lgbasallote):

welcome ^_^ and you werent off..it's a common mistake

OpenStudy (anonymous):

wel thank you!